OpenStudy (anonymous):

Calculate the discriminant. x2 – 2x + 4 = 0 A. 20 B. 0 C. 12 D. –12

OpenStudy (anonymous):

a = 1, b = -2, and c= 4 The discrminant is b^2 - 4ac. Compute b^2 - 4ac using the values of a,b, and c.

OpenStudy (anonymous):

So

OpenStudy (anonymous):

can you evaluate b^2 - 4ac when a = 1, b = -2, and c = 4?

OpenStudy (anonymous):

i got -20

OpenStudy (anonymous):

and there is no -20

OpenStudy (anonymous):

(-2)^2 - 4(1)(4) = 4 - 16 = -12

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

You added 4 to 16 to get 20. Why you did that, beats me! But at least you now understand why it is -12.